Jacob left the airport traveling toward the ocean two hours before Amanda. Amanda traveled in the opposite directions going 10 km/h slower than Jacob for two hours after which time they were 160 km apart. How fast did Jacob travel?

Found 2 solutions by Seutip, ikleyn:
Answer by Seutip(231)

Let x - be the speed of jacob in km/hr
x-10 - be the speed of amanda in km/hr
(2hours)(x)+(2hours)(x-10)+(2hours)(x)=160km
Jacob. Amanda. Jacob.
2x+2x-20+2x=160km/hr
6x-20=160
6x=180
x=30km/hr
Jacob travels 30km/hr
Amanda 20km/hr
Hope that helps! Stay safe.

Answer by ikleyn(52864)

.

The correct answer is

    Jacob travels 30 km/hr

    Amanda 20 km/hr